A parallel plate air capacitor has a capacitance of 5 F. What is the capacitance if the area of the plates is decreased to one-half its original value while keeping all other factors constant?

1
Expert's answer
2022-03-22T11:00:52-0400

The capacitance of parallel plate capacitor


"C=\frac{\epsilon_0 A}{d}""C'=\frac{\epsilon_0 A'}{d}=\frac{\epsilon_0 A/2}{d}=\frac{1}{2}\times\frac{\epsilon_0 A}{d}=\frac{1}{2}C\\
=\frac{1}{2}\times \rm 5\: F=2.5\: F"